# Transformation-free linear simplicial-simplicial regression via constrained iterative reweighted least squares

### v1

@@ Abstract
Simplicial-simplicial regression refers to the regression setting where both the responses and predictor variables lie within the simplex space, i.e. they are compositional. fiksel2022 proposed a transformation-free lienar regression model, that minimizes the Kullback-Leibler divergence from the observed to the fitted compositions was recently proposed. To effectively estimate the regression coefficients the EM algorithm was employed. We formulate the model as a constrained logistic regression, in the spirit of tsagris2025, and we estimate the regression coefficients using constrained iteratively reweighted least squares. This approach makes the estimation procedure significantly faster.
@@ 1 Introduction
Compositional data [^1] In the field of econometrics they are known as multivariate fractional data (mullahy2015, murteira2016). are non-negative multivariate vectors whose variables (typically called components) conveying only relative information. When the vectors are scaled to sum to 1, their sample space is the standard simplex
@@
Examples with both simplicial responses and predictor variables include wang2013 who modelled the linear relationship between the economic output and inputs in China, and dimarzio2015 who used a composition of the O-horizon layer to predicts composition of the moss layer. chen2017 modelled the relationship between age structure and consumption structure across economic areas. filzmoser2018 who related the percentage of males and females with different education levels across countries. ait2003, alenazi2019 examined how different methods for estimating the composition of white blood cell types are related. chen2021 examined how the the chemical metabolites of Astragali Radix and the plasma metabolites of rats after giving Astragali Radix are related. tsagris2025 studied the relationship between the crop production and the crop cultivated area in Greece and the relationship between voting proportions between successive years in Spain. Finally, rios2025 investigated the association between two high-dimensional multi-omics compositions.
@@
An issue with the EM algorithm adopted for the estimation of the TFLR’s coefficients is that the EM is in general slow. We address this issue by employing the constrained iteratively reweighted least squares (CIRLS) methodology, where the simplicial constraints are imposed on the regression coefficients. Tsagris’s implementation of the EM algorithm (tsagris2025) was shown to be 4 times faster than the implementation of fiksel2022. Our simulation studies showcase that CIRLS yields a 2-85 times speed improvement over the implementation of tsagris2025.
The next section describes the TFLR model and the EM algorithm employed to estimate the regression parameters. Section 3 delineates the CIRLS for logistic regression and adopts it to the TFLR model. Simulation studies are presented in Section 4 followed by the Conclusions section closing the paper.
@@ 2 The TFLR model
$\min\left\{\sum_{i=1}^{n}\sum_{k=1}^{Dr}y_{ik}\log\left(\frac{y_{ik}}{\sum_{j=1}^{D_{p}}B_{jk}x_{ij}}\right)\right\}=\max\left\{\sum_{i=1}^{n}\sum_{k=1}^{Dr}y_{ik}\log\left(\sum_{j=1}^{D_{p}}B_{jk}x_{ij}\right)\right\}.$ (4)
Unlike conventional approaches that rely on log-ratio transformations (ALR, CLR, ILR), their model works directly in the simplex space, avoiding transformation-induced interpretability issues. The expected value of the compositional outcome $\bm{y}$ is modeled as a single Markov transition from the compositional predictor $\bm{x}$:
$\mathbb{E}[\bm{y}|\bm{x}]=\bm{x}^{T}\bm{B}$ (5)
where $\bm{B}$ is a column-stochastic transition matrix with non-negative entries that sum to 1 in each row.
The estimation method, described below, is based on first-moment specifications using estimating equations, avoiding full likelihood specifications and providing robustness to different data-generating mechanisms. The approach naturally accommodates zeros and ones in both predictors and outcomes, addressing a major limitation of log-ratio based methods that require data pre-processing. tsagris2025 though, highlighted that the aforementioned approach allows for zero values in the response variable, but not for the simplicial predictor, in general. Think for example the case of an observation $\bm{x}_{i}=(x_{i1},x_{i2},0,0)$, where $x_{i1}$ and $x_{i2}$ are different from zero and and the $k$-th column of the estimated matrix $\bm{B}$ is equal to $(0,0,B_{3k},B_{4k})^{\top}$, where both $B_{3k}$ and $B_{4k}$ are not zero. Another possibility is when a column of the $\bm{B}$ matrix is full of zeros. These two cases will produce $x_{ik}B_{ik}=0$ and hence (4) cannot be computed. fiksel2022 bypasses this problem by adding a small quantity ($\epsilon=10^{-8}$) during computation of the KLD.
2.1 Expectation-maximisation algorithm for simplicial-simplicial regression
fiksel2022 employed an expectation-maximisation (EM) algorithm to estimate the TFLR model parameters, particularly for the column-stochastic transition matrix $\bm{B}$.
fiksel2022 framed the estimation problem as a missing data problem, where each observed compositional vector $\bm{y}_{i}$ is viewed as the result of a two-step generative process. a) For each component $j$ of $\bm{x}_{i}$, a component $k$ of $\bm{y}_{i}$ is selected with probability $b_{jk}$. b) The contributions from all components of $\bm{x}_{i}$ are aggregated according to their proportions $x_{ij}$. The steps of the EM algortihm are described below.
@@ 2.1 Expectation-maximisation algorithm for simplicial-simplicial regression
    Initialization: Set $\bm{B}^{(0)}$ to a valid column-stochastic matrix
@@
    $z_{ijk}^{(t)}=\frac{x_{ij}b_{jk}^{(t-1)}}{\sum_{j^{\prime}=1}^{p}x_{ij^{\prime}}b_{j^{\prime}k}^{(t-1)}}\cdot y_{ik}$
@@
    $\text{KLd}{(t-1)}-\text{KLD}^{(t)}<\epsilon,$
    where $\epsilon$ is a small tolerance value.
The EM updates naturally satisfy the non-negativity and sum-to-one constraints on each row of $\bm{B}$. Each iteration of the algorithm is guaranteed to increase the likelihood (or pseudo-likelihood) function. The algorithm handles numerical underflow/overflow issues by working with log-transformed probabilities when necessary. The method accommodates zeros in the data naturally (particularly in the simplicial response), without requiring pseudo-counts or other preprocessing steps.
The EM algorithm provides an elegant computational approach for this transformation-free compositional regression, ensuring that the constraints on the transition matrix are naturally maintained throughout the estimation process.
@@ 3 TFLR with CIRLS
We give some preliminaries at first regarding the TFLR, the IRLS, the CIRLS and how the latter can be used to estimate the regresion coefficients of the TFLR.
@@ 3.2 Logistic regression with identity link and simplicial constraints
where $\bm{j}_{p}=(1,\ldots,1)^{\top}$ is the p-dimensional vector of $1s$, and $\bm{I}_{p}$ is the p-dimensional identity matrix.
@@
A valid starting point for $\bm{\beta}^{(0)}$ that satisfies the simplex constraints is:
$\beta_{j}^{(0)}=\frac{1}{p}\quad\text{for }j=1,2,\ldots,p$ (9)
This gives equal weight to each coefficient. Alternatively, the starting values may come from ftting a constrained OLS with simplicial constraints imposed on the regression coefficients. The convergence criterion we employed was the absolute difference in two successive values of the log-likelihood. In some cases, the optimal solution may lie on the boundary of the simplex, with some coefficients exactly equal to 0 or 1. This can lead to degeneracy in the optimization problem. QP solvers are generally designed to handle such cases, but it may be beneficial to use a solver that employs an interior point method or another approach that handles degeneracy well.
3.3 TFLR with CIRLS
In order to formulate the TFLR as a univariate logistic regression model we follow tsagris2025 and apply the same approach. who used the vectorisation operation on the simplicial response and constructed the design matrix as a block matrix. The $\bm{D}$ matrix is a $D_{rp}\times D_{rp}$ diagonal matrix, where $D_{rp}=D_{r}\times D_{p}$, and is related to the $\bm{X}^{\top}\bm{X}$ in the following manner
@@ 4 Simulation studies
Simulation studies were conducted to compare the computational cost (running time) of the EM and the CIRLS estimation procedures for the regression coefficients of the TFLR. The simplicial response consisted of $D_{r}=(3,5,7,10)$ components, while the simplicial predictor contained $D_{p}=(5,10)$ components, and the sample sizes considered were 1,000 up to 10,000 with an increasing step size of 10,000 and from then the step size increased to 5,000 up to 50,000. The results were averaged over $100$ replicates for each combination of dimensionality and sample size.
@@ 4.1 Speed-up factors
Figure 1 visualizes the speed-up factors of CIRLS to EM. The speed-up factor increases with increasing dimensionalities, both in the simplicial responses and/or predictor variables. The EM was always slower than the CIRLs, ranging from nearly 2 times up to 50 times, when $p=5$ and from 2.5 times up to 85 times when $p=10$. Evidently, as the sample size increases the speed-up improvements decrease, mainly because there are many matrix multiplications involved, increasing the computational burden. Regarding the value of the KLD, the discrepancy between the two methods was never greater than $10^{-5}$.
@@
(a) $p=5$ (b) $p=10$
Figure 1: Speed-up factors of the two methods. The running time of EM divided by the running time of CIRLS with IRLS.

### v2

# Faster estimation of the transformation-free linear simplicial-simplicial regression via constrained iterative reweighted least squares

@@ Abstract
Simplicial-simplicial regression refers to the regression setting where both the responses and predictor variables lie within the simplex space, i.e. they are compositional. fiksel2022 proposed a transformation-free linear regression model, that minimizes the Kullback-Leibler divergence from the observed to the fitted compositions, where the EM algorithm is used to estimate the regression coefficients. We formulate the model as a constrained logistic regression, in the spirit of tsagris2025, and we estimate the regression coefficients using constrained iteratively reweighted least squares. The simulation studies depict that this algorithm makes the estimation procedure significantly faster, uses less memory, and in some cases gives a better solution.
@@ 1 Introduction
Compositional data [^1] In the field of econometrics they are known as multivariate fractional data (mullahy2015, murteira2016). are non-negative multivariate vectors whose variables (typically called components) convey only relative information. When the vectors are scaled to sum to 1, their sample space is the standard simplex
@@
Examples with both simplicial responses and predictor variables include wang2013 who modelled the linear relationship between the economic output and inputs in China, and dimarzio2015 who used the composition of the O-horizon layer to predict the composition of the moss layer. chen2017 modelled the relationship between age structure and consumption structure across economic areas. filzmoser2018 who related the percentage of males and females with different education levels across countries. ait2003, alenazi2019 examined how different methods for estimating the composition of white blood cell types are related. chen2021 examined how the the chemical metabolites of Astragali Radix and the plasma metabolites of rats after giving Astragali Radix are related. tsagris2025 studied the relationship between the crop production and the crop cultivated area in Greece and the relationship between voting proportions between successive years in Spain. Finally, rios2025 investigated the association between two high-dimensional multi-omics compositions.
@@
An issue with the EM algorithm adopted for the estimation of the TFLR’s coefficients is that the EM is in general slow. To mitigate this computational cost we employ the constrained iteratively reweighted least squares (CIRLS) algorithm, with simplicial constraints imposed on the regression coefficients. Tsagris’s implementation of the EM algorithm (tsagris2025) was shown to be 4 times faster than the implementation of fiksel2022. Our simulation studies demonstrate that CIRLS yields a 2-255 times, depending on the case, speed improvement over the implementation of tsagris2025. These speed improvements are particularly important in the following cases: a) having to analyze multiple datasets, b) performing simulation studies, c) working with high dimensional compositions (rios2025), and d) performing permutation or bootstrap based hypothesis tests (fiksel2022, tsagris2025).
The next section describes the TFLR model and the EM algorithm employed to estimate the regression parameters and Section 3 delineates the CIRLS for logistic regression and adopts it to the TFLR model. Section 4 presents the computational complexity of the both algorithms, Section 5 contains simulation studies, followed by the Conclusions section closing the paper.
@@ 2 The TFLR model
$\min\left\{\sum_{i=1}^{n}\sum_{k=1}^{Dr}y_{ik}\log\left(\frac{y_{ik}}{\sum_{j=1}^{D_{p}}x_{ij}B_{jk}}\right)\right\}=\max\left\{\sum_{i=1}^{n}\sum_{k=1}^{Dr}y_{ik}\log\left(\sum_{j=1}^{D_{p}}x_{ij}B_{jk}\right)\right\}.$ (4)
Unlike conventional approaches that rely on log-ratio transformations (ALR, CLR, ILR), the TFLR model works directly in the simplex space, avoiding transformation-induced interpretability issues. The estimation method, described below, is based on first-moment specifications using estimating equations, avoiding full likelihood specifications and providing robustness to different data-generating mechanisms. The approach naturally accommodates zeros and ones in both predictors and outcomes, addressing a major limitation of log-ratio based methods that require data pre-processing. tsagris2025 though, highlighted that the aforementioned approach allows for zero values in the response variable, but not for the simplicial predictor, in general. Consider for example the case of an observation $\bm{x}_{i}=(x_{i1},x_{i2},0,0)$, where $x_{i1}$ and $x_{i2}$ are different from zero and and the $k$-th column of the estimated matrix $\bm{B}$ is equal to $(0,0,B_{3k},B_{4k})^{\top}$, where both $B_{3k}$ and $B_{4k}$ are not zero. Another possibility is when a column of the $\bm{B}$ matrix is full of zeros. These two cases will produce $x_{ik}B_{ik}=0$ and hence (4) cannot be computed. fiksel2022 bypasses this problem by adding a small quantity ($\delta=10^{-8}$) during computation of the KLD.
2.1 Expectation-maximisation algorithm
fiksel2022 employed an expectation-maximisation (EM) algorithm to estimate the TFLR model parameters, particularly for the column-stochastic transition matrix $\bm{B}$. They framed the estimation problem as a missing data problem, where each observed compositional vector $\bm{y}_{i}$ is viewed as the result of a two-step generative process. a) For each component $j$ of $\bm{x}_{i}$, a component $k$ of $\bm{y}_{i}$ is selected with probability $b_{jk}$. b) The contributions from all components of $\bm{x}_{i}$ are aggregated according to their proportions $x_{ij}$. The steps of the EM algorithm are described below.
@@ 2.1 Expectation-maximisation algorithm
    Initialization: Set $\bm{B}^{(0)}$ to a valid column-stochastic matrix, i.e., each row is equal to $1/D_{p}$.
@@
    $z_{ijk}^{(t)}=\frac{x_{ij}B_{jk}^{(t-1)}}{\sum_{j^{\prime}=1}^{p}x_{ij}B_{jk}^{(t-1)}}\cdot y_{ik}$
@@
    $\|\bm{B}^{(t)}-\bm{B}^{(t-1)}\|_{1}<\epsilon\ \ \text{or}\ \ \text{KLd}^{(t-1)}-\text{KLD}^{(t)}<\epsilon,$
    where $\|.\|_{1}$ denotes the $L_{1}$ norm and $\epsilon$ is a small tolerance value.
The EM updates naturally satisfy the non-negativity and sum-to-one constraints on each row of $\bm{B}$. Each iteration of the algorithm is guaranteed to increase the likelihood (or pseudo-likelihood) function. The algorithm handles numerical underflow/overflow issues by working with log-transformed probabilities when necessary and accommodates zeros in the data naturally (particularly in the simplicial response), without requiring zero value imputation.
In order to avoid for() loops, by vectorization, in the EM algorithm, the implementation in the R package Compositional (compositional2025) constructs the following design matrix
$\widetilde{\bm{X}}=\bm{I}_{D_{r}}\otimes\bm{X}=\left(\begin{array}[]{cccc}\bm{X}&\bm{0}&\ldots&\bm{0}\\ \bm{0}&\ddots&\bm{0}&\vdots\\ \vdots&\bm{0}&\ddots&\bm{0}\\ \bm{0}&\ldots&\bm{0}&\bm{X}\end{array}\right).$ (9)
This resulted in an implementation that was shown to be around 4 times faster (tsagris2025) than the implementation in the R package codalm (codalm2021), at the cost of adding extra memory requirements.
@@ 3 TFLR with CIRLS
We give some preliminaries at first regarding the TFLR, the IRLS, the CIRLS and how the latter can be used to estimate the regression coefficients of the TFLR.
@@ 3.2 Logistic regression with identity link and simplicial constraints
where $\bm{j}_{p}=(1,\ldots,1)^{\top}$ is the p-dimensional vector of $1s$, and $\bm{I}_{p}$ is the p-dimensional identity matrix. The first row of $\bm{A}$ corresponds to the sum of the $\beta_{s}$ that must be equal to 1, which is denoted by the first element of the $\bm{b}$ vector. Hence the $p+1$ constraints, are in fact one equality constraint and $p$ inequality constraints.
@@
A valid starting point for $\bm{\beta}^{(0)}$ that satisfies the simplex constraints could be
$\beta_{j}^{(0)}=\frac{1}{p}\quad\text{for }j=1,2,\ldots,p.$
This gives equal weight to each coefficient. Alternatively, the starting values may come from fitting a constrained OLS with simplicial constraints imposed on the regression coefficients. The convergence criterion we employed was the absolute difference in two successive values of the log-likelihood. In some cases, the optimal solution may lie on the boundary of the simplex, with some coefficients exactly equal to 0 or 1. This can lead to degeneracy in the optimization problem. QP solvers are generally designed to handle such cases, but it may be beneficial to use a solver that employs an interior point method or another approach that handles degeneracy well.
3.3 Customizing CIRLS for TFLR
In order to formulate the TFLR as a univariate logistic regression model we follow tsagris2025 and apply the same approach. They constructed the $\bm{D}$ matrix is a $D_{rp}\times D_{rp}$ diagonal matrix, where $D_{rp}=D_{r}\times D_{p}$, and is related to the $\bm{X}^{\top}\bm{X}$ in the following manner
@@ 3.3 Customizing CIRLS for TFLR
Effectively, we do not vectorise the simplicial responses, or the compositional predictors, and evidently neither the weighting matrix $\bm{W}$. We obtain the $\bm{B}$ matrix from the QP, and subsequently we compute the matrix with the fitted responses, and then the KLD value. We then update the weighting matrix and this process iterates until convergence of the KLD. The initial matrix $\bm{B}$ is obtained from the constrained least squares method of tsagris2025. The dual active-set method of goldfarb1983, implemented in the R package quadprog (quadprog2019), is used during the QP minimization.
4 Computational complexity of EM and CIRLS algorithms
We analyze the computational and memory complexity of both the EM and CIRLS algorithms to understand their scalability properties and explain the observed speed-up factors.
4.1 Per-iteration Time Complexity
  • EM Algorithm: The E-step computes the latent allocations $z_{ijk}^{(t)}$ for all $i\in n$, $j\in D_{p}$, and $k\in D_{r}$. For each triplet $(i,j,k)$, the E-step costs $\mathcal{O}(nD_{p}^{2}D_{r})$. The M-step performs normalization in $\mathcal{O}(nD_{p}D_{r})$ time. Thus, the complexity of the $t$-th iteration is
    $T_{\text{EM}}^{(t)}=\mathcal{O}(nD_{p}^{2}D_{r}).$
  • CIRLS Algorithm: After vectorization, the main computational steps are: a) computing the linear predictor $\bm{\mu}^{(t)}=\tilde{\mathbf{X}}\tilde{\bm{\beta}}^{(t)}$ in $\mathcal{O}(nD_{p}D_{r})$ time; b) computing the weight matrix $\mathbf{W}^{(t)}$ in $\mathcal{O}(nD_{r})$ time; c) forming $\mathbf{D}=\tilde{\mathbf{X}}^{\top}\mathbf{W}^{(t)}\tilde{\mathbf{X}}$, which is block diagonal with $D_{r}$ blocks of size $D_{p}\times D_{p}$, requiring $\mathcal{O}(nD_{p}^{2}D_{r})$ operations; and d) solving the QP problem with $D_{rp}=D_{r}D_{p}$ variables and $\mathcal{O}(D_{rp})$ constraints. Using interior-point methods, the QP solve costs $\mathcal{O}(D_{rp}^{3})=\mathcal{O}(D_{p}^{3}D_{r}^{3})$. The complexity of the $t$-th iteration is
    $T_{\text{CIRLS}}^{(t)}=\mathcal{O}(nD_{p}^{2}D_{r}+D_{p}^{3}D_{r}^{3}).$
For typical problems where $n\gg D_{p}D_{r}$, both algorithms have similar per-iteration complexity $\mathcal{O}(nD_{p}^{2}D_{r})$, as the matrix operations dominate the QP overhead. In this regime, the speed-up of CIRLS over EM arises primarily from requiring fewer iterations to converge (typically $K_{\text{CIRLS}}\approx\frac{1}{2}K_{\text{EM}}$ to $\frac{1}{5}K_{\text{EM}}$). For problems with large $D_{p}\times D_{r}$ relative to $n$, the QP solving becomes the computational bottleneck in CIRLS. However, this can be mitigated by exploiting the fact that the $\mathbf{B}$ matrix has $D_{r}$ independent rows (each satisfying simplex constraints), allowing decomposition into $D_{r}$ separate QP problems of size $D_{p}$ that can be solved in parallel.
If EM requires $K_{\text{EM}}$ iterations and CIRLS requires $K_{\text{CIRLS}}$ iterations to converge, the total complexities are
$T_{\text{EM}}$ $=\mathcal{O}(K_{\text{EM}}\cdot nD_{p}^{2}D_{r}),$
$T_{\text{CIRLS}}$ $=\mathcal{O}(K_{\text{CIRLS}}\cdot[nD_{p}^{2}D_{r}+D_{p}^{3}D_{r}^{3}]).$
4.2 Memory Complexity
  • EM Algorithm: The main memory requirements are: a) the data matrices $\mathbf{X}$ and $\mathbf{Y}$, requiring $\mathcal{O}(n(D_{p}+D_{r}))$ space; b) the coefficient matrix $\mathbf{B}$, requiring $\mathcal{O}(D_{p}D_{r})$ space; and c) the latent allocation array $z_{ijk}$, requiring $\mathcal{O}(nD_{p}D_{r})$ space. Thus,
    $M_{\text{EM}}=\mathcal{O}(nD_{p}D_{r}).$
  • CIRLS Algorithm: In addition to the data and coefficients, CIRLS requires: a) the block-diagonal matrix $\mathbf{D}$, which can be stored efficiently as $D_{r}$ blocks of size $D_{p}\times D_{p}$, requiring $\mathcal{O}(D_{p}^{2}D_{r})$ space; b) the QP solver workspace, typically requiring $\mathcal{O}(D_{p}^{2}D_{r}^{2})$ space for storing constraint matrices and factorizations. Thus,
    $M_{\text{CIRLS}}=\mathcal{O}(nD_{p}D_{r}+D_{p}^{2}D_{r}^{2}).$
[…]
Table 1: Computational and memory complexity comparison of EM and CIRLS algorithms.
This is shown later in the simulation studies.
EM CIRLS
Per-iteration time $\mathcal{O}(nD_{p}^{2}D_{r})$ $\mathcal{O}(nD_{p}^{2}D_{r}+D_{p}^{3}D_{r}^{3})$
Iterations to converge $K_{\text{EM}}$ $K_{\text{CIRLS}}<K_{\text{EM}}$
Total time $\mathcal{O}(K_{\text{EM}}nD_{p}^{2}D_{r})$ $\mathcal{O}(K_{\text{CIRLS}}[nD_{p}^{2}D_{r}+D_{p}^{3}D_{r}^{3}])$
Memory $\mathcal{O}(nD_{p}D_{r})$ $\mathcal{O}(nD_{p}D_{r}+D_{p}^{2}D_{r}^{2})$
4.3 Implementation details in R
Table 1 presents the theoretical complexities of the two algorithms. However, the R implementations are slightly different. which is of dimensions $\left(n\times D_{rp}\right)\times\left(n\times D_{rp}\right)$. As stated earlier, the EM algorithm’s R implementation in the package Compositional (compositional2025), constructs the design matrix $\widetilde{\bm{X}}$ of Eq. (9). This save running time but increases the memory requirements to $\mathcal{O}(nD_{p}^{2}D_{r})$. As for CIRLS algorithm, its bottleneck is the computation of the block matrix $\bm{D}$ (17) that involves performing $2\times D_{r}$ cross-product matrix multiplications.
Second, computational optimizations have been applied. Instead of computing the full KLD, at each iteration of either algorithm, only the parts involving the fitted values are computed, i.e. the right hand side of Eq. (4) is computed. The matrix $\bm{D}$ (17), required by QP, is initialised once, and its elements are updated at every iteration.
Both algorithms are available in the R package Compositional (compositional2025). The relevant functions are tflr() and tflr.irls() and they return the same output, the running time, the number of iterations required, the KLD value, the estimated regression coefficients and optionally, the fitted values for new observations.
@@ 5 Simulation studies
Simulation studies [^4] The experiments were conducted utilising a Dell laptop equipped with Intel Core i5-1053G1 CPU at 1GHz, with 256 GB SSD, 8 GB RAM and Windows installed. were conducted to compare the computational cost (running time) of the EM and the CIRLS estimation procedures for the regression coefficients of the TFLR. For all cases, both simplicial responses and predictors were generated from a Dirichlet distribution. The sample sizes considered were 1,000 up to 10,000 with an increasing step size of 10,000 and from then the step size increased to 5,000 up to 50,000. For all cases, the tolerance value to check convergence was $\epsilon=10^{-6}$, and the results were averaged over $100$ replicates for each combination of dimensionality and sample size. Two scenarios were considered: a) the responses and the predictor variables are independent and b) they are linearly dependent, with the following configurations:
  • The simplicial response consisted of $D_{r}=(3,5,7,10)$ components, while the simplicial predictor contained $D_{p}=(5,10)$ components.
  • The simplicial response consisted of $D_{r}=(20,30,40,50)$ components, while the simplicial predictor contained $D_{p}=(15,20)$ components.
EM-1 refers to the implementation in the package Compositional, while EM-2 refers to the implementation of (fiksel2022) in the package codalm.
@@ 5.1 Speed-up factors
Figure 1 visualizes the speed-up factors of CIRLS to EM when the two compositions are independent. When $p=5$, the speed-up factors of CIRLS compared to EM-1 range from 1.3 up to 50, whereas compared to EM-2 they range from 8 up to 50. When $p=10$, the EM-1 is 1.6 up 220 times slower, whereas EM-2 is 12-141 times slower. With $p=15$ and $p=20$, EM-1 becomes extremely slow, slower than EM-2, due to large memory requirements and the speed-up factors are not presented. EM-2 is 17-50 times slower than CIRLS when $p=15$, and 22-64 times slower than CIRLS when $p=20$. Regarding the KLD, both algorithms returned the same value.
Figure 2 visualizes the speed-up factors when the compositions are dependent. When $p=5$, the speed-up factors of CIRLS compared to EM-1 range from 4 up to 227, whereas compared to EM-2 they range from 2 up to 96. When $p=10$, the EM-1 is 9 up 347 times slower, whereas EM-2 is 12-264 times slower. With $p=15$ and $p=20$, EM-1 becomes extremely slow, slower than EM-2, due to large memory requirements and the speed-up factors are not presented. EM-2 is 46-114 times slower than CIRLS when $p=15$, and 40-157 times slower than CIRLS when $p=20$. Notably, there was a discrpeancy in the produced KLD values. CIRLS always produced slightly higher KLD than the EM algorithm. The discrepancy was always in the 4rth, the 5th or the 6th digit.
@@
(a) EM-1 to CIRLS when $p=5$ (b) EM-2 to CIRLS when $p=5$
Refer to caption Refer to caption
(c) EM-1 to CIRLS when $p=10$ (d) EM-2 to CIRLS when $p=10$
Refer to caption Refer to caption
(e) EM-2 to CIRLS when $p=15$ (f) EM-2 to CIRLS when $p=20$
Figure 1: Speed-up factors of the two algorithms, when the two compositions are independent: running time of EM divided by the running time of CIRLS with IRLS. The horizontal golden line indicates when the unity speed-up factor, when the running times are equal.
Refer to caption Refer to caption
(a) EM-1 to CIRLS when $p=5$ (b) EM-2 to CIRLS when $p=5$
Refer to caption Refer to caption
(c) EM-1 to CIRLS when $p=10$ (d) EM-2 to CIRLS when $p=10$
Refer to caption Refer to caption
(e) EM-2 to CIRLS when $p=15$ (f) EM-2 to CIRLS when $p=20$
Figure 2: Speed-up factors of the two algorithms, when the two compositions are dependent: running time of EM divided by the running time of CIRLS with IRLS. The horizontal golden line indicates when the unity speed-up factor, when the running times are equal.